OpenStudy (anonymous):

Perform the operations: (5+2i)-(8-3i)

OpenStudy (anonymous):

so far I have the following: -40-15i+16i-6i^2

OpenStudy (anonymous):

@Hero

hero (hero):

(5 + 2i)-(8 - 3i) = (5 - 8) + (2i + 3i) = -3 + 5i Remember it's (5 + 2i) MINUS (8 - 3i) NOT (5 + 2i) TIMES (8 - 3i)

OpenStudy (anonymous):

so you dont distribute

hero (hero):

No, you SUBTRACT

hero (hero):

The minus between the two binomials actually means something believe it or not.

OpenStudy (anonymous):

I see that I was looking at an example that did not have the subtraction in the middle

OpenStudy (anonymous):

what gets me also is the answer -3+5i

hero (hero):

That is actually the right answer.

OpenStudy (anonymous):

ok I see that thanks for your help